OpenStudy (anonymous):

How do i write a quadratic equation in standard form with the given roots 1,4

OpenStudy (anonymous):

I would work backwards from factoring.

OpenStudy (anonymous):

Start with x=1 and x=4 x-1=0 and x-4=0 (x-1)(x-4)=0 Then FOIL

OpenStudy (anonymous):

so are the 1 and 4 negative or positive?

OpenStudy (anonymous):

The roots you gave are positive.

OpenStudy (anonymous):

so would it be \[x ^{2}+4x+1x+4=0 \]

OpenStudy (anonymous):

Almost. Check your negatives.

OpenStudy (anonymous):

What ones are negatives? I have trouble with understanding these things.

OpenStudy (anonymous):

Remember that the roots are the numbers that are equal to x. So a root of 1 is written as x=1. Then set it equal to 0 by subtracting 1 from both sides giving you x-1=0